# On the convergence of bootstrap current to the Shaing-Callen limit in stellarators

### v1

@@ 5 Effect of particle precession
It should be noted that $\langle v_{g}^{\vartheta_{0}}\rangle_{b}$ here should be evaluated over the real bounce orbit with finite Larmor radius, which is in contrast to $\langle v_{g}^{r}\rangle_{b}$ where this results only in a $\rho_{L}$ correction to Eq. (111). Such an evaluation is required to retain in $\langle v_{g}^{\vartheta_{0}}\rangle_{b}$ the effect of magnetic shear which is of the order one if $\Omega_{E}\lesssim\Omega_{B}$ (see (Shaing, 2015; Albert et al., 2016; Martitsch et al., 2014)). An alternative way is to use Eqs. (3.45) of Morozov and Solov’ev (1966) resulting in $\langle v_{g}^{r}\rangle_{b}=J_{b}^{-1}\partial J_{\parallel}/\partial\vartheta_{0}$ and $\langle v_{g}^{\vartheta_{0}}\rangle_{b}=-J_{b}^{-1}\partial J_{\parallel}/\partial r$ where $J_{\parallel}=m\oint{\rm d}lv_{\parallel}$ is the parallel adiabatic invariant computed along the field line, which makes the Liouville’s theorem $\partial(J_{b}\langle v_{g}^{r}\rangle_{b})/\partial r+\partial(J_{b}\langle v_{g}^{\vartheta_{0}}\rangle_{b})/\partial\vartheta_{0}=0$ obvious.
@@
Eq. (135) means that off-set effect is generally non-local, i.e. radial transport due to entrapping of passing particles at some given surface is produced in some radial range $\delta r_{c}$ (radial correlation length) which in the $1/\nu$ regime is of the order $\delta r_{c}\sim\tau_{d}\langle v_{g}^{r}\rangle_{b}$, where $\tau_{d}\sim\varepsilon_{M}/\nu$ is the de-trapping time, and, at lower collisionalities, where $\langle v_{g}^{\vartheta_{0}}\rangle_{b}>1/\tau_{d}$ and mild electric fields, $\Omega_{E}\leq\Omega_{B}$, is of the size of radial deviation of precessing trapped orbit which does not scale with Larmor radius and is fully determined by the magnetic field geometry. Thus, local ansatz (1) has a limited validity in the latter case where transport coefficients and bootstrap current need a non-local treatment, e.g., by Monte-Carlo methods (Sasinowski and Boozer, 1995; Satake et al., 2008) (an exception is weakly perturbed tokamak equilibria where a quasi-local ansatz (Martitsch et al., 2014) can be used). Of course, $\delta r_{c}$ must be small compared to the plasma radius for the magnetic fields intended for the reactor, however, it is not necessarily small compared to the radial variation scale of the off-set which is quite sensitive to small modulations of the field and $\iota$ changes. One consequence of this non-locality is discussed in more details in Section 6.
@@
Eq. (135) should be solved as is if the off-set is produced in the wide poloidal range containing the trapped-passing boundary layer, which is the case of almost aligned maxima discussed in Sections 3.3, 3.5 and 4.3. However, in a more typical case with a distinct global maximum (or two global maxima, which are also possible in stellarator symmetry), the width of the off-set region $\delta\vartheta$ is small (this is a typical width of a single strap in Figs. 19-22). Therefore, convective term with $\langle v_{g}^{\vartheta_{0}}\rangle_{b}$ dominates over $\langle v_{g}^{r}\rangle_{b}$ which, therefore, can be ignored in Eq. (135) together with the dependence on $\vartheta_{0}$ of $\langle v_{g}^{\vartheta_{0}}\rangle_{b}\equiv\Omega^{\vartheta}$, $J_{b}$ and $D^{J_{\perp}J_{\perp}}$, which should be taken at the global maximum point $\vartheta_{0}=\vartheta_{\max}$, thus reducing Eq. (135) to
@@
For further estimates, we assume a fast precession case, $\Omega^{\vartheta}\tau_{d}\gg\delta\vartheta$, which would be the condition of the $\sqrt{\nu}$ transport regime (Galeev and Sagdeev, 1979) in case $\delta\vartheta\sim 1$ but is established here for smaller precession frequencies due to $\delta\vartheta\ll 1$. In this case, solution is localized in $\eta$ near the matching boundary $\eta_{m}$, and we can ignore class transitions which occur in the well trapped particle domain and also set $\eta=\eta_{m}$ in all the coefficients. Introducing the dimensionless variables as follows,
@@
$D_{31}^{\ast}=D_{31}^{\rm SC}+\left(D_{31}^{1/\nu}-D_{\rm 31}^{\rm SC}\right)\left(\frac{\nu_{\ast}E_{\rm ref}}{\nu_{\ast}^{\rm ref}E_{r}}\right)^{1/2},$ (143)
where $D_{31}^{\rm SC}$ is Shaing-Callen asymptotic value, $D_{31}^{1/\nu}$ is the actual value in the $1/\nu$ regime, and $\nu_{\ast}^{\rm ref}$ and $E_{\rm ref}$ are the values of $\nu_{\ast}$ and $E_{r}$ for a single reference point in the plot. As one can see in Fig. 27, scaling (143) well represents the results with mild electric field but fails for two largest values of $v_{E}^{\ast}$ where the (tokamak like) boundary layer off-set (Helander et al., 2011) becomes important (see also Fig. 17) and where the condition $\Omega_{E}\tau_{b}/\delta\vartheta\ll 1$ becomes violated.
@@ 6 Off-set in the direct problem, bootstrap effect at the magnetic axis
For the equilibrium (Maxwellian) particle distribution which is constant on the flux surface, detrapping process is balanced by entrapping of passing particles leading to zero phase space fluxes (detailed equilibrium) and, respectively, to zero parallel current. In the presence of radial gradient, the distribution function of trapped particles is not constant on the flux surface anymore. It is larger in the region with positive bounce averaged radial drift, and smaller where it is negative. In stellarator symmetry, it is obvious that particles are detrapped in those regions in opposite directions, thus leading to the overall shift of passing particle distribution and, respectively, to the parallel current. Destroyed symmetry of the ripple wells is described by the source term $s_{(1)}$, Eq. (4), which generates “anti-particles” in the wells with $\langle v_{g}^{r}\rangle_{b}>0$ and “particles” where $\langle v_{g}^{r}\rangle_{b}<0$. In the asymptotical $1/\nu$ regime, these particles and anti-particles produced by the source term $s_{(1)}$ within ripple wells can leave these wells only through boundary layers between classes, where they are re-distributed and annihilated. As we saw in Sections 2.2 and 2.3, parallel flows within these boundary layers make a contribution of the order one to the equilibrium Pfirsch-Schlüter current. This changes at finite collisionality where “relevant ripples” appear with both limiting local maxima entering the trapped-passing boundary layer. Re-distribution of particles and anti-particles between “relevant” ripples occurs then through the trapped passing boundary layer (this layer is empty in the asymptotical $1/\nu$ regime) causing the parallel flow there and respective asymmetry of the distribution function over the pitch parameter $\lambda$. Such an odd distribution in the boundary layer, $g_{\text{odd}}$, serves as a boundary condition which shifts from zero the odd part of the distribution function in the whole passing particle domain thus generating a bootstrap current.
We can estimate $g_{\text{odd}}$ in the boundary layer using the argument at the end of Section 2.2 as $g_{\text{odd}}\sim H_{j}(\eta_{\text{loc}})/\delta\eta_{\rm ref}$ where $H_{j}(\eta_{\text{loc}})$ is the collisional flux through class boundary $\eta=\eta_{\text{loc}}\equiv\max(\eta_{j},\eta_{j+1})$ of the relevant ripple (which corresponds to the off-set domain in the adjoint problem) with
@@ 7 Conclusion
For the analysis of boundary layer effects and resulting off-set of bootstrap current we have developed a simplified propagator method, and obtained with its help, besides the condition of equivalent ripples (89) formulated earlier in (Helander et al., 2011) as a property of ideal quasi-isodynamic stellarators leading there to a tokamak like bootstrap effect, also the expression (108) for the off-set of the distribution function in case of imperfectly aligned local maxima and slightly violated ripple equivalence condition (108) together with respective estimate (134) of the resulting off-set of bootstrap coefficient in case of the devices with non-optimized bounce-averaged radial drift of trapped particles. Actually, the case of nearly aligned maxima is the most interesting from the point of view of good fusion alpha particle confinement. This is naturally the case in quasi-symmetric configurations and in quasi-isodynamic configurations with poloidally closed of contours of $B$. As it can be seen from condition (134), the way to avoid the off-set of bootstrap current in such configurations is to align the maxima accurately and to realize ripple equivalence condition (89). The latter condition is almost the same with the condition of tied to the flux surface contours of parallel adiabatic invariant $J_{\parallel}$ for barely passing particles which, once realized in case of perfectly aligned maxima, would prevent collisionless trapped-passing particle transitions and resulting stochastic transport. The essential difference of the respective integral $I_{j}$ from $J_{\parallel}=m\oint{\rm d}l\,v_{\parallel}$ is only in an extra $1/B$ factor in the sub-integrand. Both conditions are naturally satisfied if quasi-symmetry conditions are realized exactly which, however, cannot be achieved at all flux surfaces (Garren and Boozer, 1991). More generally, these conditions are fulfilled simultaneously for perfectly omnigeneous fields (Cary and Shasharina, 1997; Helander and Nührenberg, 2009) being also an idealization (Cary and Shasharina, 1997).
@@
As follows from Section 4.3, the strongest bootstrap current off-set due to aligned maxima appears for the above parameters, $\varepsilon_{M}\sim\varepsilon_{t}$, $\iota=1$ and $N_{\text{tor}}=5$ if the mis-match of maxima is smaller than $\Delta B/B\sim 2.5\%$ for the middle of the collisionality range (smaller then $8\%$ for the highest collisionality). According to Eq. (134), bringing this off-set back to the level of bootstrap current in the equivalent tokamak would require to align the maxima better than $\Delta B/B\sim 0.3\%$ for the lowest collisionality and to fulfill ripple equivalence condition to within $\Delta A_{o}\lesssim 50\%$. The latter condition does not appear to be too restrictive and can probably be achieved already when tying to the flux surface $J_{\parallel}$ contours for barely passing particles.
@@
Unfortunately, the off-set in the distribution function cannot be further minimized by better fulfilling the above conditions because there remains a residual off-set containing the term $\propto\log\nu_{\ast}$ not accounted for in this paper. Nevertheless, the resulting bootstrap current off-set can still be minimized by reducing the integral bounce-averaged drift represented by factors $I_{\cal V}$, Eq. (114), which are well defined in case of aligned maxima by magnetic field geometry alone. Certain minimization of these factors can be achieved at mild collisionalities where off-set wells are relatively short already as a by-product of minimization of $1/\nu$ transport coefficients (29) via $\varepsilon_{\text{eff}}$. This appears to be the case in W-7X “standard” configuration where the off-set in $D_{31}^{\ast}$ shown in Fig. 3 of (Kernbichler et al., 2016) is of the order of Shaing-Callen value (10 % of the equivalent tokamak value) for collisionalities $\nu^{\ast}>10^{-5}$, and where $\varepsilon_{\text{eff}}$ is reduced by about an order of magnitude compared to a standard stellarator (see Fig. 10 in (Beidler et al., 2011)). Strong $D_{31}^{\ast}$ off-set of the order one shows up there only at very low collisionalities where it is due to the (non-optimized) bounce-averaged drift of long bananas (off-set wells are long).
@@
Of course, the above simple recipes assuming independent minimization of purely geometrical quantities (asymptotic $\lambda_{bB}$ and $I_{\cal V}$) do not apply in the general case without strong alignment of field maxima where multiple off-set well types contribute simultaneously with their relative contributions depending on collisionality and precession velocity. Minimization of bootstrap current in such cases requires direct numerical evaluation of all transport coefficients and the resulting plasma parameter profiles at finite plasma collisionality (Geiger et al., 2015). For the evaluation of bootstrap coefficient in such cases, computations of $\lambda_{bB}$ in the $1/\nu$ regime would be useful if one accounts for the precession by bounce-averaged approach outlined in Section 5. A particular tool for this is NEO-2 which has been specially designed for the effective evaluation of bootstrap current.

### v3

@@ 5 Effect of particle precession
It should be noted that due to the use of field aligned (Clebsch-like) spatial variables both, $\langle v_{g}^{r}\rangle_{b}$ and $\langle v_{g}^{\vartheta_{0}}\rangle_{b}$ can be evaluated here along the field lines (see (Shaing, 2015; Calvo et al., 2017; d’Herbemont et al., 2022)). On the contrary, for the non-aligned variables (periodic flux coordinates) tangential components of the bounce-averaged drift should be evaluated over the real bounce orbit with finite Larmor radius, which is in contrast to $\langle v_{g}^{r}\rangle_{b}$ where this results only in a $\rho_{L}$ correction to Eq. (108). Such an evaluation is required to retain in the resulting precession frequency the effect of magnetic shear which is of the order one if $\Omega_{E}\lesssim\Omega_{B}$ (see (Shaing, 2015; Albert et al., 2016; Martitsch et al., 2016)). A way equivalent to bounce averaging velocity components is to use Eqs. (3.45) of Morozov and Solov’ev (1966) resulting in $\langle v_{g}^{r}\rangle_{b}=J_{b}^{-1}\partial J_{\parallel}/\partial\vartheta_{0}$ and $\langle v_{g}^{\vartheta_{0}}\rangle_{b}=-J_{b}^{-1}\partial J_{\parallel}/\partial r$ where $J_{\parallel}=m\oint{\rm d}lv_{\parallel}$ is the parallel adiabatic invariant computed along the field line, which makes the Liouville’s theorem $\partial(J_{b}\langle v_{g}^{r}\rangle_{b})/\partial r+\partial(J_{b}\langle v_{g}^{\vartheta_{0}}\rangle_{b})/\partial\vartheta_{0}=0$ obvious.
@@
Eq. (133) means that off-set effect is generally non-local, i.e. radial transport due to entrapping of passing particles at some given surface is produced in some radial range $\delta r_{c}$ (radial correlation length) which in the $1/\nu$ regime is of the order $\delta r_{c}\sim\tau_{d}\langle v_{g}^{r}\rangle_{b}$, where $\tau_{d}\sim\varepsilon_{M}/\nu$ is the de-trapping time, and, at lower collisionalities, where $\langle v_{g}^{\vartheta_{0}}\rangle_{b}>1/\tau_{d}$ and mild electric fields, $\Omega_{E}\leq\Omega_{B}$, is of the size of radial deviation of precessing trapped orbit which does not scale with Larmor radius and is fully determined by the magnetic field geometry. Thus, local ansatz (1) has a limited validity in the latter case where transport coefficients and bootstrap current need a non-local treatment, e.g., by Monte-Carlo methods (Sasinowski and Boozer, 1995; Satake et al., 2008) (an exception is weakly perturbed tokamak equilibria where a quasi-local ansatz (Martitsch et al., 2016) can be used retaining the boundary layer effects). Of course, $\delta r_{c}$ must be small compared to the plasma radius for the magnetic fields intended for the reactor, however, it is not necessarily small compared to the radial variation scale of the off-set which is quite sensitive to small modulations of the field and $\iota$ changes. One consequence of this non-locality is discussed in more details in Section 6.
@@
Eq. (133) should be solved as is if the off-set is produced in the wide poloidal range containing the trapped-passing boundary layer, which is the case of almost aligned maxima discussed in Sections 3.3, 3.5 and 4.3. However, in a more typical case with a distinct global maximum (or two global maxima, which are also possible in stellarator symmetry), the width of the off-set region $\delta\vartheta$ is small (this is a typical width of a single strap in Figs. 21-24). Therefore, convective term with $\langle v_{g}^{\vartheta_{0}}\rangle_{b}$ dominates over $\langle v_{g}^{r}\rangle_{b}$ which, therefore, can be ignored in Eq. (133) together with the dependence on $\vartheta_{0}$ of $\langle v_{g}^{\vartheta_{0}}\rangle_{b}\equiv\Omega^{\vartheta}$, $J_{b}$ and $D^{J_{\perp}J_{\perp}}$, which should be taken at the global maximum point $\vartheta_{0}=\vartheta_{\max}$, thus reducing Eq. (133) to
@@
For further estimates, we assume a fast precession case, $\Omega^{\vartheta}\tau_{d}\gg\delta\vartheta$ (with $\Omega^{\vartheta}\tau_{b}\ll\delta\vartheta$ staying fulfilled, what is possible at low collisionalities where $\tau_{b}\ll\tau_{d}$), which would be the condition of the $\sqrt{\nu}$ transport regime (Galeev and Sagdeev, 1979) in case $\delta\vartheta\sim 1$ but is established here for smaller precession frequencies due to $\delta\vartheta\ll 1$. In this case, solution is localized in $\eta$ near the matching boundary $\eta_{m}$, and we can ignore class transitions which occur in the well trapped particle domain and also set $\eta=\eta_{m}$ in all the coefficients. Introducing the dimensionless variables as follows,
@@
$D^{\ast}_{31}=D_{31}^{\rm SC}+\left(D_{31}^{1/\nu}-D_{\rm 31}^{\rm SC}\right)\frac{D^{\ast}_{\text{ref}}-D_{\rm 31}^{\rm SC}}{D_{\text{ref}}^{1/\nu}-D_{\rm 31}^{\rm SC}}\left(\frac{\nu_{\ast}E_{\rm ref}}{\nu_{\rm ref}E_{r}}\right)^{1/2},$ (141)
where $D_{31}^{\rm SC}$ is the Shaing-Callen asymptotic value of $D^{\ast}_{31}$, $D_{31}^{1/\nu}$ is the actual value in the $1/\nu$ regime, $\nu_{\rm ref}$ and $E_{\rm ref}$ are the values of $\nu_{\ast}$ and $E_{r}$ for a single reference point in the plot, and $D_{\text{ref}}$ and $D_{\text{ref}}^{1/\nu}$ are the values of $D^{\ast}_{31}$ for $(\nu_{\ast},E_{r})=(\nu_{\rm ref},E_{\rm ref})$ and $(\nu_{\ast},E_{r})=(\nu_{\rm ref},0)$, respectively. As one can see in Fig. 31, scaling (141) well represents the results with mild electric field but fails for two largest values of $v_{E}^{\ast}$ where the (tokamak like) boundary layer off-set (Helander et al., 2011) becomes important (see also Fig. 19) and where the condition $\Omega_{E}\tau_{b}/\delta\vartheta\ll 1$ becomes violated.
@@ 6 Off-set in the direct problem, bootstrap effect at the magnetic axis
For the equilibrium (Maxwellian) particle distribution which is constant on the flux surface, detrapping process is balanced by entrapping of passing particles leading to zero phase space fluxes (detailed equilibrium) and, respectively, to zero parallel current. In the presence of radial gradient, the distribution function of trapped particles is not constant on the flux surface anymore. It is larger in the region with positive bounce averaged radial drift, and smaller where it is negative. In stellarator symmetry, it is obvious that particles are detrapped in those regions in opposite directions, thus leading to the overall shift of passing particle distribution and, respectively, to the parallel current. Destroyed symmetry of the ripple wells is described by the source term $s_{(1)}$, Eq. (4), which generates “anti-particles” in the wells with $\langle v_{g}^{r}\rangle_{b}>0$ and “particles” where $\langle v_{g}^{r}\rangle_{b}<0$. In the asymptotical $1/\nu$ regime, these particles and anti-particles produced by the source term $s_{(1)}$ within ripple wells can leave these wells only through boundary layers between classes, where they are re-distributed and annihilated. As we saw in Sections 2.2 and 2.3, parallel flows within these boundary layers make a contribution of the order one to the equilibrium Pfirsch-Schlüter current. This changes at finite collisionality where “relevant ripples” appear with both limiting local maxima entering the trapped-passing boundary layer. Re-distribution of particles and anti-particles between “relevant” ripples occurs then through the trapped passing boundary layer (this layer is empty in the asymptotical $1/\nu$ regime) causing the parallel flow there and respective asymmetry of the distribution function over the pitch parameter $\lambda$. Such an odd distribution in the boundary layer, $g^{\text{odd}}$, serves as a boundary condition which shifts from zero the odd part of the distribution function in the whole passing particle domain thus generating a bootstrap current.
We can estimate $g^{\text{odd}}$ in the boundary layer using the argument at the end of Section 2.2 as $g_{\text{odd}}\sim H_{j}(\eta_{\text{loc}})/\delta\eta_{\rm ref}$ where $H_{j}(\eta_{\text{loc}})$ is the collisional flux through class boundary $\eta=\eta_{\text{loc}}\equiv\max(\eta_{j},\eta_{j+1})$ of the relevant ripple (which corresponds to the off-set domain in the adjoint problem) with
@@
Note that we have retained in the estimates (143) and (144) only the off-set effect but ignored the usual bootstrap mechanism where radial correlation length $\tau_{cj}$ corresponds to the banana width which is much smaller for electrons than $\tau_{cj}$ resulting from the $1/\nu$ transport. However, this usual mechanism is not small for the ions whose radial orbit extent and, respectively, radial correlation length is much larger (with the orbits near the axis comprising a variety of classes even in an axisymmetric tokamak (Shaing et al., 2015; Buchholz et al., 2022)), and, more importantly, who retain due to their high mass nearly all the momentum within the species during Coulomb collisions. For the axisymmetric devices, this mechanism generates the parallel ion flow (generally different from the $E\times B$ flow) but not the current since this flow is fully compensated by the electron flow driven by the collisional friction. However, if the toroidal ripple is present, it brakes the electrons leading to some steady current on axis if the ion flow is finite there. The latter is affected by ripple too via the neoclassical toroidal viscous torque (Shaing, 2015; Albert et al., 2016; Martitsch et al., 2016) which tends to bring the toroidal rotation velocity towards the (finite) intrinsic velocity driven by the ion temperature gradient. Due to the finite radial orbit width, finite gradients at surrounding flux surfaces result in finite ion rotation on axis which finally leads to a finite equilibrium current there. Generally, once the axial symmetry of the tokamak magnetic field is violated, there appear various mechanisms resulting in finite non-inductive toroidal current on axis which is necessary for the steady state tokamak operation, see, e.g., (Weening and Boozer, 1992). At the same time, mechanisms discussed here are rather weak for that purpose which is better achieved with help of stellarator-tokamak hybrids (Ku and Boozer, 2009; Henneberg and Plunk, 2024; Liang et al., 2025).
@@ 7 Conclusion
For the analysis of boundary layer effects and resulting off-set of bootstrap current we have developed a simplified propagator method, and obtained with its help, besides the condition of equivalent ripples (85) formulated earlier in (Helander et al., 2011) as a property of ideal quasi-isodynamic stellarators leading there to a tokamak like bootstrap effect, also the expression (105) for the off-set of the distribution function in case of imperfectly aligned local maxima and slightly violated ripple equivalence condition (105) together with respective estimate (132) of the resulting off-set of bootstrap coefficient in case of the devices with non-optimized bounce-averaged radial drift of trapped particles. Actually, the case of nearly aligned maxima is the most interesting from the point of view of good fusion alpha particle confinement. This is naturally the case in quasi-symmetric configurations and in quasi-isodynamic configurations with poloidally closed of contours of $B$. As it can be seen from condition (132), the way to avoid the off-set of bootstrap current in such configurations is to align the maxima accurately and to realize ripple equivalence condition (85). The latter condition is almost the same as the condition of being tied to the flux surface contours of the parallel adiabatic invariant $J_{\parallel}$ for barely passing particles which, once realized in case of perfectly aligned maxima, would prevent collisionless trapped-passing particle transitions and resulting stochastic transport. The essential difference of the respective integral $I_{j}$ from $J_{\parallel}=m\oint{\rm d}l\,v_{\parallel}$ is only in an extra $1/B$ factor in the sub-integrand. Both conditions are naturally satisfied if quasi-symmetry conditions are realized exactly, which, however, cannot be achieved at all flux surfaces (Garren and Boozer, 1991). More generally, these conditions are fulfilled simultaneously for perfectly omnigeneous fields (Cary and Shasharina, 1997; Helander and Nührenberg, 2009) being also an idealization (Cary and Shasharina, 1997).
@@
As follows from Section 4.3, the strongest bootstrap current off-set due to nearly aligned maxima appears for the above parameters, $\varepsilon_{M}\sim\varepsilon_{t}$, $\iota=1$ and $N_{\text{tor}}=5$ if the mis-match of maxima is around $\Delta B/B\sim 2.5\%$ for the middle of the collisionality range (around $8\%$ for the highest collisionality). According to Eq. (132), bringing this off-set back to the level of bootstrap current in the equivalent tokamak would require to align the maxima better than $\Delta B/B\sim 0.3\%$ for the lowest collisionality and to fulfill ripple equivalence condition to within $\Delta A_{o}\lesssim 50\%$. The latter condition does not appear to be too restrictive and can probably be achieved already when tying to the flux surface $J_{\parallel}$ contours for barely passing particles.
@@
Unfortunately, the off-set in the distribution function cannot be further minimized by better fulfilling the above conditions because there remains a residual off-set containing the term $\propto\log\nu_{\ast}$ not accounted for in this paper. Nevertheless, the resulting bootstrap current off-set can still be minimized by reducing the integral bounce-averaged drift represented by factors $I_{\cal V}$, Eq. (111), which are well-defined in case of aligned maxima by magnetic field geometry alone. Certain minimization of these factors can be achieved at mild collisionalities where off-set wells are relatively short already as a by-product of minimization of $1/\nu$ transport coefficients (29) via $\varepsilon_{\text{eff}}$. This appears to be the case in W-7X “standard” configuration where the off-set in $D_{31}^{\ast}$ shown in Fig. 3 of (Kernbichler et al., 2016) is of the order of Shaing-Callen value (10 % of the equivalent tokamak value) for collisionalities $\nu^{\ast}>10^{-5}$, and where $\varepsilon_{\text{eff}}$ is reduced by about an order of magnitude compared to a standard stellarator (see Fig. 10 in (Beidler et al., 2011)). Strong $D_{31}^{\ast}$ off-set of the order one shows up there only at very low collisionalities where it is due to the (non-optimized) bounce-averaged drift of long bananas (off-set wells are long).
@@
Of course, the above simple recipes assuming independent minimization of purely geometrical quantities (asymptotic $\lambda_{bB}$ and $I_{\cal V}$) do not apply in the general case without strong alignment of field maxima where multiple off-set well types contribute simultaneously with their relative contributions depending on collisionality and precession velocity. Minimization of bootstrap current in such cases requires direct numerical evaluation of all transport coefficients and the resulting plasma parameter profiles at finite plasma collisionality (Geiger et al., 2015). For the evaluation of bootstrap coefficient in such cases, computations of $\lambda_{bB}$ in the $1/\nu$ regime would be useful if one accounts for the precession by bounce-averaged approach outlined in Section 5. A particular tool for this is NEO-2 which has been specially designed for the effective evaluation of bootstrap current. Another option would be a fast neoclassical code MONKES (Escoto et al., 2024, 2025) which, similarly to NEO-2, has an efficiently parallelizable algorithm.